A 1000kg automobile moving with a speed of 24 m/s collides with a 500kg car initially at rest. If the two stick together, what is the velocity (in m/s) of the two cars after the collision relative to an automobile moving in the same direction at 15 m/s?
Answer #1

momentum before collision = momentum after collision

m1*v1 + m2*v2 = (m1+m2)*V

(1000*24)+(50*0)= (1000+500)*V


V = 16 m/s
Answer #3

Solution After collisi on thefinal velocity is samefor both From conservation of mom entum 1000kg (24m/s)0 (1000kg +500kg)v v